# Balanced Mixed-Type Tabular Data Synthesis with Diffusion Models

### v1

@@ 6 Conclusion
The objective of this study is to make progress in the area of fair machine learning. Tabular datasets sometimes contain inherent bias, such as imbalanced distributions in sensitive attributes. Training machine learning models on biased datasets may result in decisions that could negatively affect minority groups. By mitigating biases present in tabular datasets through the generation of equitable synthetic data, our approach contributes to fostering equitable decision-making processes across industries such as finance, healthcare, and employment. Moreover, in instances where sharing datasets becomes necessary, the utilization of fair synthetic data ensures the preservation of user privacy and avoids causing emotional distress to minority groups. An adverse possibility is that bad individuals could manipulate distributions in sensitive attributes to generate biased (even stronger than original) data to harm minority groups.
@@ A.1.1 UCI Adult
The UCI Adult dataset (Kohavi et al., 1996) is widely used as a benchmark for exploring fairness and bias in machine learning. It contains 48842 data points. Each data point has 14 attributes and a binary target variable indicating whether an individual earns over fifty thousand dollars annually. The dataset encompasses employment, education, and demographic information, and sensitive features are sex and race. Download it from OpenML.
A.1.2 Bank Marketing
@@
The Bank Marketing dataset is collected from direct marketing campaigns of a Portuguese banking institution. Comprising 45211 instances with 16 features, the predicted outcome is a binary variable indicating whether a client subscribed to a term deposit. Demographic, economic, and past marketing campaign data are included in features, and marital status is sensitive. Download it from OpenML.
A.1.3 Cardio
The Cardio dataset focuses on diagnosing cardiovascular disease (yes or no) and it contains 70000 instances. It incorporates 11 features, which contain demographics and factual information (blood pressure, cholesterol levels, etc.), as well as self-reported labels such as whether smoking or not. The sensitive feature is sex. Download it from OpenML.
A.1.4 Credit Card
The Credit Card dataset (Yeh & Lien, 2009) explores customer default payments (yes or no) among credit card clients in Taiwan, and it contains 30000 data points. Each data point has 23 features including the information of demographics and financial records (monthly payments, etc.). The sensitive features are sex and education. Download it from OpenML.
A.1.5 Depression
The Depression dataset contains demographic and clinical data (concentration, rumination, etc.) to predict a diagnosis of depression. It has 11000 instances, each with 18 features, and sensitive features are sex and race. Download it from OpenML.
A.1.6 KDD Census
The KDD Census dataset is an expanded version of the UCI Adult dataset. It involves 199523 instances, each featuring 41 attributes and a binary target variable indicating income over fifty thousand dollars per year. The dataset covers employment, education, and demographic details, with sex and race as sensitive features. Download it from OpenML.
A.1.7 Law School
The Law School dataset presents a survey on the law school admission process. It has 20800 instances. Each instance has 12 attributes and a binary target variable indicating admission to law school. Demographic and academic information, such as GPA, are included in features. The sensitive features are sex and race. Download it from OpenML.
A.2 Number of Features
Dataset Number of Numerical Features Number of Categorical Features
UCI Adult 6 8
Bank Marketing 7 9
Cardio 5 6
Credit Card 14 9
Depression 14 4
KDD Census 7 34
Law School 6 6
Table 6: The number of numerical attributes and categorical attributes in experimental datasets
A.3 Evaluation of Best-Performing Classifiers on Original Datasets
Dataset Model Choice Accuracy Demographic Parity Ratio Equalized Odds Ratio
UCI Adult XGBoost $0.866$ $0.353$ $0.273$
Bank Marketing XGBoost $0.900$ $0.619$ $0.540$
Cardio XGBoost $0.733$ $0.992$ $0.989$
Credit Card Random Forest $0.816$ $0.778$ $0.834$
Depression XGBoost $0.941$ $0.696$ $0.644$
KDD Census XGBoost $0.956$ $0.151$ $0.170$
Law School XGBoost $1.000$ $0.825$ $1.000$
Table 7: The values of accuracy and fairness scores of the best-performing classifier on experimental datasets
A.4 Class Imbalance
(a) Bank Marketing
(b) Cardio
(c) Credit Card
(d) Depression
(e) KDD Census
(f) Law School
Figure 4: Visualizations of class imbalance in sensitive attributes in experimental datasets except UCI Adult

### v3

@@ 7 Conclusion
The objective of this study is to make progress in the area of fair machine learning. Tabular datasets sometimes contain inherent bias, such as imbalanced distributions in sensitive attributes. Training machine learning models on biased datasets may result in decisions that could negatively affect minority groups. By mitigating biases present in tabular datasets through the generation of equitable synthetic data, our approach contributes to fostering equitable decision-making processes across industries such as finance, healthcare, and employment. Moreover, in instances where sharing datasets becomes necessary, the utilization of fair synthetic data ensures the preservation of user privacy and avoids causing emotional distress to minority groups. An adverse possibility is that bad individuals could manipulate distributions in sensitive attributes to generate biased (even stronger than original) data to harm minority groups. Additionally, the widespread release of synthetic data can diminish the quality of data sources available on the internet. Repeatedly training generative models on synthetic data in a self-consuming loop can lead to a decline in the quality of data synthesis, as discussed in (Alemohammad et al., 2023).
@@ A.1 Adult
The Adult dataset (Kohavi et al., 1996) is widely used as a benchmark for exploring fairness and bias in machine learning. It contains 48842 data points. Each data point has 14 attributes and a binary target variable indicating whether an individual earns over fifty thousand dollars annually. The dataset encompasses employment, education, and demographic information, and sensitive features are sex and race. Download it from OpenML.
A.2 Bank Marketing
@@
The Bank Marketing dataset (Moro et al., 2014) is collected from direct marketing campaigns of a Portuguese banking institution. Comprising 45211 instances with 16 features, the predicted outcome is a binary variable indicating whether a client subscribed to a term deposit. Demographic, economic, and past marketing campaign data are included in features, and marital status is sensitive. Download it from OpenML.
A.3 COMPAS
The COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) dataset (Larson et al., 2016) is derived from a risk assessment tool used in the criminal justice system to evaluate the likelihood of recidivism among criminal defendants. It includes data on individuals arrested in Broward County, Florida, with 8 features and a binary outcome indicating whether the individual was predicted to reoffend. Features encompass demographic information, criminal history, and COMPAS risk scores, with sex and race being sensitive attributes. Download it from OpenML.
Appendix B Reproducibility
We performed our experiments on Ubuntu 20.04.2 LTS, utilizing Python version 3.10.14. Our framework of choice was PyTorch 2.3.0, with CUDA version 12.1, and we ran our computations on an NVIDIA GeForce RTX 3090. The hyperparameters we used for our method are listed as follows. The variable names can be found in our attached code. We tune the hyperparameters using Optuna Akiba et al. (2019) to optimize the validation AUC.
Table 5: The values of hyperparameters for our method.
Notation Variable Meaning Value/Search Space
- batch_size Batch size $256$
$\delta$ warmup_steps Warm-up timestep 00
$\omega_{s}$ cond_guid_weight Sensitive guidance weight $1.0$
$\lambda$ cond_guid_threshold Threshold of the “security gate“ $1.0$
$\omega_{m}$ cond_momentum_weight Momentum weight $0.5$
$\beta$ cond_momentum_beta Momentum correction factor $0.5$
$\omega_{g}$ overall_guid_weight Guidance weight $1.0$
$T$ n_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
- n_epochs Number of training epochs $\{100,500,1000\}$
- lr Learning rate $[0.00001,0.003]$
For deep learning-based methods, we use the original hyperparameter settings provided in their respective implementations and mainly fine-tune the number of epochs, learning rate, and number of diffusion timesteps (if applicable) to optimize performance. For SMOTE-based methods, we adjust the number of nearest neighbors to achieve better results. We have detailed the hyperparameter search spaces for all baseline methods in Table 6. The hyperparameter search spaces are the same for all datasets.
Table 6: The values of hyperparameters for baseline methods.
Method Variable Meaning Value/Search Space
CoDi total_epochs_both Number of training epochs $\{100,500,1000,3000\}$
n_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
GReaT batch_size Batch size $\{4,8\}$
n_epochs Number of training epochs $\{5,10,20\}$
SMOTE knn Number of nearest neighbors for interpolation $\{2,21\}$
STaSy lr Learning rate $[0.00001,0.003]$
n_epochs Number of training epochs $\{100,500,1000\}$
TabDDPM lr Learning rate $[0.00001,0.003]$
n_epochs Number of training epochs $\{100,500,1000\}$
num_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
TabSyn lr Learning rate $[0.001,0.002]$
n_epochs Number of training epochs $\{4000\}$
FairCB knn Number of nearest neighbors for interpolation $\{2,21\}$
FairTGAN lr Learning rate $[0.00001,0.003]$
n_epochs Number of training epochs $\{100,500,1000\}$
fair_epochs Number of fairness-related training epochs $\{100,500\}$
Appendix C Efficiency
Table 7 provides a summary of the number of parameters, training time, and sampling time for various tabular data synthesis methods. Our proposed method is moderately sized, with 482,545 parameters, which is larger than TabDDPM but significantly smaller than most deep learning-based models, except FairTGAN. In terms of training efficiency, our method achieves a training time of 11.8 seconds per epoch, which is comparable to CoDi (10.5 seconds per epoch). However, the sampling time is a significant limitation, as it takes approximately 1,200 seconds to generate a synthetic dataset of the same size as the training data, making it substantially slower than all other baseline methods.
Table 7: Comparison of training and sampling efficiency across methods
Method Number of Parameters Training Time (s/epoch) Sampling Time (s)
CoDi 2206711 10.5 9.1
GReaT 81912576 725.5 236.3
SMOTE - 51.1 17.0
STaSy 10653326 1.4 25.1
TabDDPM 109672 6.0 118.9
TabSyn 10575912 0.6 2.7
FairCB - 64.3 5.4
FairTGAN 45368 1.9 0.3
Ours 482545 11.8 1213.9
Appendix D Stacked Bar Plots of Contingency Tables
D.1 Adult
Figure 5: Comparison of models on the Adult dataset.
D.2 Bank Marketing
Figure 6: Comparison of models on the Bank Marketing dataset.
D.3 COMPAS
Figure 7: Comparison of models on the COMPAS dataset.
Appendix E Additional Numerical Results
E.1 Performance and Fairness Metrics with Sensitive Features Replaced by Uniform Distribution
We present performance and fairness of CatBoost trained on datasets with sensitive attribute replaced by uniform distribution in Table 8 and Table 9 respectively.
Table 8: The values of machine learning efficiency with the best classifier for each dataset. The sensitive attributes are replaced by uniform distribution.
Adult Bank COMPAS Mean
[…]
Real .$928_{\pm.000}.$$936_{\pm.000}.$$799_{\pm.000}.$ $88.8\%$
CoDi .$858_{\pm.002}.$$827_{\pm.015}.$$678_{\pm.008}.$ $78.8\%$
GReaT .$900_{\pm.003}.$$690_{\pm.026}.$$714_{\pm.010}.$ $76.8\%$
SMOTE .$914_{\pm.000}.$$928_{\pm.001}.$$772_{\pm.002}.$ $87.1\%$
STaSy .$883_{\pm.000}.$$894_{\pm.003}.$$722_{\pm.014}.$ $83.3\%$
TabDDPM .$906_{\pm.001}.$$917_{\pm.002}.$$745_{\pm.002}.$ $85.6\%$
TabSyn .$910_{\pm.000}.$$918_{\pm.001}.$$745_{\pm.001}.$ $85.8\%$
FairCB .$914_{\pm.000}.$$907_{\pm.001}.$$765_{\pm.001}.$ $86.2\%$
FairTGAN .$885_{\pm.002}.$$873_{\pm.006}.$$703_{\pm.002}.$ $82.0\%$
Ours .$895_{\pm.001}.$$913_{\pm.002}.$$732_{\pm.001}.$ $84.7\%$
Table 9: The values of DPR and EOR with best classifier for each dataset. The sensitive attributes are replaced by uniform distribution.
Adult Bank COMPAS Mean Adult Bank COMPAS Mean
[…]
Real .$307_{\pm.004}.$$415_{\pm.008}.$$790_{\pm.005}.$ $50.4\%$$192_{\pm.008}.$$368_{\pm.008}.$$857_{\pm.007}.$ $47.2\%$
CoDi .$345_{\pm.030}.$$402_{\pm.063}.$$901_{\pm.011}.$ $54.9\%$$328_{\pm.038}.$$556_{\pm.125}.$$923_{\pm.010}.$ $60.2\%$
GReaT .$256_{\pm.010}.$$622_{\pm.269}.$$814_{\pm.024}.$ $56.4\%$$169_{\pm.017}.$$405_{\pm.079}.$$802_{\pm.055}.$ $45.9\%$
SMOTE .$331_{\pm.017}.$$413_{\pm.025}.$$777_{\pm.028}.$ $50.7\%$$276_{\pm.045}.$$359_{\pm.026}.$$826_{\pm.041}.$ $48.7\%$
STaSy .$304_{\pm.072}.$$532_{\pm.060}.$$780_{\pm.072}.$ $53.9\%$$240_{\pm.106}.$$542_{\pm.065}.$$739_{\pm.073}.$ $50.7\%$
TabDDPM .$281_{\pm.004}.$$341_{\pm.017}.$$750_{\pm.031}.$ $45.7\%$$191_{\pm.010}.$$339_{\pm.010}.$$816_{\pm.040}.$ $44.9\%$
TabSyn .$286_{\pm.005}.$$319_{\pm.018}.$$815_{\pm.016}.$ $47.3\%$$182_{\pm.011}.$$285_{\pm.022}.$$847_{\pm.016}.$ $43.8\%$
FairCB .$305_{\pm.005}.$$623_{\pm.011}.$$762_{\pm.023}.$ $56.3\%$$227_{\pm.006}.$$680_{\pm.021}.$$824_{\pm.023}.$ $57.7\%$
FairTGAN .$481_{\pm.021}.$$812_{\pm.115}.$$692_{\pm.026}.$ $66.2\%$$544_{\pm.036}.$$567_{\pm.169}.$$732_{\pm.024}.$ $61.4\%$
Ours .$503_{\pm.017}.$$690_{\pm.085}.$$781_{\pm.006}.$ $65.8\%$$580_{\pm.027}.$$631_{\pm.087}.$$781_{\pm.015}.$ $66.4\%$